### v1

Abstract
The price of oil can rise because of a disruption to supply or an increase in demand. The nature of the price change determines the dynamic effects. As Kilian (2009a) put it: “not all oil price shocks are alike.” Using the latest available data, we extend Kilian’s analysis using the R ecosystem and provide more evidence for Kilian’s conclusions. Inference based on unknown conditional heteroskedasticity strengthens the conclusions. With the updated shocks, we assess how a local economy responds to the global oil market, an application that is relevant to policymakers concerned with the transition away from fossil fuels.
Keywords: Kern, local labor market, oil price, real economic activity, structural vector autoregression, unemployment rate, vector autoregression
JEL Codes: E24, E31, E32, Q41, Q43, Q48, R23
1 Introduction
In a stylized model for the oil market, the price of oil depends on supply and demand. Price can go up if demand shifts outward or supply shifts inward. Even though the nature of the shocks differs, each causes the price of oil to rise. Surprisingly, as Kilian (2009a) points out, much theoretical, empirical, and policy-oriented work focuses on the effects of oil-price fluctuations, ignoring the nature of shocks. This risks flawed policy. Fluctuations in the price of oil are an amalgam of demand and supply shocks. The dynamics generated by a demand shock differ from the dynamics generated by a supply shock. At any point in time, the composition of shocks may be far from the average, making correlations between the price of oil and US macroeconomic aggregates potentially meaningless and uninformative for policymakers (Kilian, 2008, 2009a).
In brief, here is the main issue: Kilian (2009a) establishes that the dynamic effects of an increase in the price of oil depend on the nature of the increase—whether the increase was the result of a disruption to supply or an increase in demand. The conclusions are based on a structural model of the global market for oil estimated over the period February 1973 to December 2007. Do the same conclusions hold using all available data? Are the results robust to inference that allows for conditional heteroskedasticity?
Using data from the US Energy Information Administration on oil production and price and data available through FRED on global real economic activity, we estimate the structural vector autoregressive model of the global crude-oil market proposed by Kilian (2009a). The updated results provide further evidence that Kilian’s (2009a) conclusions hold. In addition, a procedure that accounts for conditional heteroskedasticity of unknown form in the VAR model, indicates that Kilian’s conclusions are even stronger than initially stated (Brüggemann, Jentsch, and Trenkler, 2016).
The structural VAR model disentangles supply shocks from demand shocks in the global crude-oil market. In the model, supply comprises global oil production and demand comprises two components—a component that reflects the global business cycle and a component that reflects oil-specific demand (buyers concerned about availability in the future will buy out of precaution, for example). This allows Kilian (2009a) to show how each shock separately determines fluctuations in the real price of crude oil. Moreover, how real Gross Domestic Product responds to an unanticipated increase in the price of oil depends on the nature of the shock. The same conclusion applies to the Consumer Price Index. Kilian’s results explain the macroeconomic implications of oil-price fluctuations.
In a different direction, we use the updated series of structural shocks to learn about another important policy question: how does a local labor market respond to the identified structural shocks? The question is of interest to policymakers in California. California has adopted the goal of carbon neutrality by 2045 or earlier (CARB, 2022). In Kern County, which accounts for 70 percent of California’s oil production, there is concern about potential job loss and lost tax revenue as the economy transitions away from fossil fuels. Yet, how much Kern’s economy depends on oil is hard to assess. We document that the unemployment rate in Kern responds meaningfully to identified shocks in the global market for oil and that the response depends on the nature of the shock. This is suggestive evidence that policymakers may need to consider expanding options in Kern through, for example, place-based policy. [^1] What will happen in California as 2045 approaches is impossible to know with certainty. In particular, a backward-looking structural VAR model will not be a good predictor of abrupt change. Instead, we view Kern’s responses to identified structural innovations as suggestive evidence that the Kern economy depends on oil. The exercise shows how Kilian’s (2009a) pioneering methods and conclusions apply more generally.
Questions about how oil affects the US economy are not fully answered. Further research will benefit from other contributions of our replication. We port Kilian’s original MATLAB code to R. Some researchers may prefer to work in R as the R ecosystem makes tidying data easier. We use the targets package to ensure replicability. The targets package provides Make-like tools that automatically make any necessary changes to dependencies when any change is registered. In addition, we retrieve data from the US Energy Information Administration and FRED using code, which makes estimating the most recent structural shocks as straightforward as rerunning the code. These resources will make addressing questions easier in the ongoing debate about the effects oil-price shocks have on the US economy. [^2] An introduction to Make is provided by Jesús Fernández-Villaverde in lecture notes titled “Computational Methods for Economists.”
We present a replication of Kilian’s (2009a) results in section 2. We conclude that Kilian (2009a) should be praised for writing and sharing code that makes reproducing results straightforward (Kilian, 2009b). The replication comprises a description of the data (section 2.1), a specification of the structural VAR model (section 2.2), a presentation of the estimated impulse response functions (section 2.3), a decomposition of the real price of oil into identified components (section 2.4), and a second-stage analysis of how US real GDP and the Consumer Price Index respond to the identified structural shocks. In section 3 we strengthen Kilian’s (2009a) conclusions by estimating the structural VAR model using data from February 1974 to March 2024 and constructing confidence intervals for impulse responses that are robust to arbitrary forms of conditional heteroskedasticity. To demonstrate the wider applicability of the results, in section 4 we analyze how the global oil market affects outcomes in a local labor market.
2 Replication
2.1 Data
Separating supply shocks from demand shocks requires data on oil production and the price of oil. Data on production are available from the US Energy Information Administration (EIA). [^3] Using the link in the text, the data can be found under “Petroleum and Other Liquids” and “Crude oil including lease condensate.” Data on price are also available from the EIA under the series U.S. Crude Oil Composite Acquisition Cost by Refiners. The EIA website includes links for retrieving the data through a web application programming interface or API. Through the API, it is possible to send structured HTTP requests that return data with JSON format or JavaScript Object Notation format. The code that gathers our data uses the EIA’s API to retrieve the latest data for production and price.
The time series that measures the global business cycle is the Index of Global Real Economic Activity, which is available through FRED. [^4] Federal Reserve Bank of Dallas, Index of Global Real Economic Activity [IGREA], retrieved from FRED, Federal Reserve Bank of St. Louis; https://fred.stlouisfed.org/series/IGREA. The series “is derived from a panel of dollar-denominated global bulk dry cargo shipping rates and may be viewed as a proxy for the volume of shipping in global industrial commodity markets.” Kilian (2009a) initially proposed this measure of the global business cycle. Kilian and Zhou (2018, 55) discuss the merits of the series, concluding “that the Kilian (2009a) index and indicies based on real commodity prices have clear advantages in modeling industrial commodity markets compared with alternative indicators of global real economic activity, including proxies for global industrial production.”
Figure 1 depicts the three series used in the structural VAR model. All series are monthly. The top panel depicts the change in global crude-oil production. The solid, blue line shows the data used by Kilian (2009a), and the dashed, red line shows data retrieved using the code for this paper (see Kilian, 2009b, for the original data and code). The original series begin in February 1973 and go through December 2007. The updated series begin in February 1974 and go through March 2024. The discrepancy occurs because data on production begin in 1974. Kilian (2009a) extends the series through 1973 using a procedure described by Barsky and Kilian (2001), a procedure that is not undertaken here. The two series largely coincide. [^5] As can be confirmed in the code, the original production series was multiplied by 10 in order to produce the plot. The reason for this transformation is unclear. Scaling any series will have no meaningful effect on the structural VAR model. Nevertheless, it speaks to the aims of the Journal of Comments and Replications in Economics, which emphasize code that reproduces results from original data sources.
Figure 1: Determinants of the global market for crude oil
Note: Top, change in global crude-oil production; middle, global real economic activity measured using the index proposed by Kilian (2009a); bottom, real price of oil measured by the acquisition cost of refiners and deflated using the Consumer Price Index for All Urban Consumers. The solid, blue lines show the data originally used by Kilian (2009a). The dashed, red lines show the updated data. All data are monthly. The updated data begin in February 1974 and go through March 2024. The original data begin one year earlier.
Sources: US Energy Information Administration, Federal Reserve Bank of Dallas, and US Bureau of Labor Statistics
The middle panel in figure 1 shows the time series of global real economic activity. Again, the solid, blue line depicts the original data and the dashed, red line depicts the updated data. A few years before 2005, the two series diverge meaningfully. This discrepancy has to do with a coding error in the original analysis described by Kilian (2019).
The bottom panel in figure 1 shows the time series for the real price of crude oil. Here the difference between the original and updated series is unknown to us. Nevertheless, the correlation between the two series is 0.98.
As Kilian (2009a) demonstrates, the three time series in figure 1 are key determinants of the global market for crude oil.
2.2 The structural VAR model
To be precise about how the determinants of the global market for crude oil are related, consider a VAR model for $y_{t}=\left(\Delta\text{oil production},\text{real activity},\text{real price of oil}\right)^{\prime}$. These are the three series shown in figure 1. The structural VAR representation is
$B_{0}y_{t}=\beta+B_{1}y_{t-1}+\cdots+B_{24}y_{t-24}+\varepsilon_{t},$ (1)
where the $B_{i}$ are $3\times 3$ parameter matricies and $\varepsilon_{t}\sim\left(0,\Sigma_{\varepsilon}\right)$; that is, $\varepsilon_{t}$ denotes a vector of serially and mutually uncorrelated innovations. The covariance matrix $\Sigma_{\varepsilon}$ is diagonal. The corresponding reduced-form model is
$y_{t}=\alpha+A_{1}y_{t-1}+\cdots+A_{24}y_{t-24}+\epsilon_{t},$ (2)
where $A_{i}=B_{0}^{-1}B_{i}$ and $\epsilon_{t}=B_{0}^{-1}\varepsilon_{t}\sim\left(0,\Sigma_{\epsilon}\right)$ is a martingale-difference sequence with positive-definite covariance matrix $\Sigma_{\epsilon}=\left(B_{0}^{-1}\right)\Sigma_{\varepsilon}\left(B_{0}^{-1}\right)^{\prime}$. Going between the structural model in (1) and the reduced-form model in (2) requires an estimate of $B_{0}^{-1}$. Short-run identifying restrictions make estimating this term straightforward.
The recursive structure for $B_{0}^{-1}$ implies that the reduced-form error $\epsilon_{t}$ can be decomposed as
$\epsilon_{t}\equiv\begin{bmatrix}[l]\epsilon_{t}^{\Delta\text{oil production}}\\ \epsilon_{t}^{\text{real activity}}\\ \epsilon_{t}^{\text{real price of oil}}\end{bmatrix}=\begin{bmatrix}b_{11}&0&0\\ b_{21}&b_{22}&0\\ b_{31}&b_{32}&b_{33}\end{bmatrix}\begin{bmatrix}[l]\varepsilon_{t}^{\text{oil supply}}\\ \varepsilon_{t}^{\text{aggregate demand}}\\ \varepsilon_{t}^{\text{oil-specific demand}}\end{bmatrix}.$ (3)
Within the system of three equations, the first equation indicates that the stochastic short-run supply curve is vertical. In other words, oil production does not respond to innovations in demand within the month. The maintained assumption is that decisions about oil supply are based on medium-run expectations and, because adjusting supply is costly, short-run demand innovations are postulated to have a negligible affect on medium-run expectations. Evidently, the model allows energy prices to respond to all past information. There are two demand shocks and each is identified by delay restrictions. The short-run demand curve is downward sloping and can be shifted by innovations to aggregate demand and innovations to oil-specific demand (such as precautionary-demand innovations). The delay restriction is that oil-specific demand shocks do not affect global real economic activity as measured by fluctuations in shipping rates relative to trend. These postulations are further discussed by Kilian (2008) and Kilian (2009a).
In (1), $B_{0}^{-1}$ is lower triangular. A lower-triangular Cholesky decomposition can therefore be used to decompose the estimate for $\widehat{\Sigma}_{\epsilon}$, starting from the reduced-form estimates of the model in (2). Uncovering the structural shocks is possible using the estimated relationships in (3).
Inference is based on the wild bootstrap discussed by Gonçalves and Kilian (2004) with 1,000 replications. Kilian (2009a) uses 2,000 replications but the results are similar. The wild bootstrap allows for conditional heteroskedasticity of unknown form. But the asymptotic validity of the wild bootstrap is for the reduced-form impulse responses and not the asymptotic distribution of the error–covariance matrix. This feature is shown by Brüggemann, Jentsch, and Trenkler (2016) and discussed by Kilian and Lütkepohl (2017). In section 3, we adopt the inference procedures proposed by Brüggemann, Jentsch, and Trenkler (2016) to investigate whether the larger confidence intervals matter in practice.
2.3 How global oil production, real economic activity, and the real price of oil respond to demand and supply shocks in the crude-oil market
The structural impulse responses define the responses of elements of $y_{t}$ to one-time impulses in $\varepsilon_{t}$. Figure 2 shows how global oil production, real economic activity, and the real price of oil respond to one-standard deviation structural innovations. The innovations are normalized so that each innovation will tend to raise the price of oil or, put differently and slightly imprecisely, figure 2 shows negative supply shocks and positive demand shocks. Figure 2 is comparable to figure 3 in Kilian (2009a).
Figure 2: Responses to one-standard-deviation structural shocks (based on original data)
Note: Solid lines show the point estimates and the shaded regions show two-standard-error bands. Estimates are based on the structural VAR model in (1), using the original data (Kilian, 2009b) (February 1973–December 2007). Confidence intervals were constructed using a recursive-design wild bootstrap. Dotted lines show the point estimates and two-standard-error bands produced by Kilian’s (2009a) code. The striped effect indicates the two sets of point estimates agree.
Figure 2 uses data provided by Kilian (2009b). These are the data shown in figure 1 with solid, blue lines. The data start in February 1973 and end in December 2007. The goal of figure 2 is to assess whether the R code written for the replication yields similar results to the MATLAB code that produces the original graphics.
Figure 2 confirms that the two sets of code produce similar results. The solid, blue lines show the point estimates of the structural impulse responses. They match those produced by the code provided by Kilian (2009a) exactly. Kilian’s (2009a) original estimates are shown using red, dotted lines. These overlay the solid, blue lines, which produces a striped effect. The shaded areas depict the two-standard-error bands constructed by the wild bootstrap. The dotted, red lines show those produced by the original code. Again, there is overwhelming overlap, which gives confidence that the two sets of codes are working.
Figure 3 uses our code base and our data over the original sample period. These data are the dashed, red lines in figure 1. There are two differences to keep in mind when comparing figure 2 to figure 3. First, the updated data on production start in February 1974. Second, the data on global real economic activity have been corrected (Kilian, 2019). Notwithstanding these differences, the patterns are strikingly similar.
Figure 3: Responses of one-standard-deviation structural shocks (based on updated data)
Note: Solid lines show the point estimates and the shaded regions show one- and two-standard-error bands. Estimates are based on the structural VAR model in (1), using the updated data (February 1974–December 2007). Confidence intervals were constructed using a recursive-design wild bootstrap.
An unexpected disruption to oil supply causes an immediate decline in oil production. The decline is partially reversed within the first 5 months, but “this pattern is consistent with the view that oil supply contractions in one region tend to trigger production increases elsewhere in the world” (Kilian, 2009a, 1061–1062). The fall in production causes a rise in the price of oil, but the rise is not overwhelmingly statistically significant. In figure 3, the disruption has little effect on real economic activity, which differs slightly from Kilian’s (2009a) observation that a supply disruption “causes a small temporary reduction of real economic activity that is partially statistically significant” (1,062). (The assessment is based on one-standard-error bands, which are not shown here. Figure 3, constructed with all available data, shows one-standard-error bands.)
Turning to the second row of figure 3, the effect of an unanticipated rise in aggregate demand causes oil production to rise about 6 months after the shock. The expansion in oil production lasts until 15 months after the shock. The unanticipated rise in aggregate demand causes a significant rise in global real economic activity. The effect is persistent. Expansions in real economic activity cause a significant rise in the real price of oil. The real price of oil rises on impact (although not statistically significantly) and the rise continues for a year.
The third row of figure 3 shows the effect of an unanticipated rise in oil-specific demand. An unanticipated rise causes oil production to briefly fall, but over the 15-month horizon, oil-specific-demand increases do not cause oil production to change much. Real economic activity in response to the unanticipated rise in oil-specific demand expands—albeit modestly, if at all. What is most clear, though, is that the unanticipated rise in oil-specific demand causes the real price of oil to jump upward immediately and continue to rise for a few months, which is consistent with models of precautionary demand. The rise in the real price of oil is highly statistically significant.
The results in 3, which use the replication code and updated data, are entirely consistent with the results in Kilian (2009a).
2.4 The cumulative effect of oil demand and oil supply shocks on the real price of oil
How much each structural shock explains of the historically observed price of crude oil can be assessed through a historical decomposition of the covariance-stationary VAR model. The historical decomposition decomposes $y_{t}$ into structural innovations through its moving-average representation. It attributes the cumulative effect of the structural innovations in oil production, real economic activity, and oil-specific demand to the real price of oil at each point in time.
Figure 4 shows the cumulative contribution of oil-supply, aggregate-demand, and oil-specific-demand shocks to the real price of oil in separate panels. Each panel shows 4 lines. The top panel shows the cumulative effect of oil-supply shocks. The solid, purple line is generated from the original code and original data. The dashed, beige line is generated by using the replication code together with the original data. The two series are numerically equal, which can be seen by the striped pattern. The dot–dashed, orange line is generated by using the replication code and the updated data through December 2007. The dotted, pink line is generated by using the replication code and updated data through March 2024.
Figure 4: Historical decomposition of the real price of oil
Note: Top, how much oil-supply shocks explain of the historically observed fluctuations in the real price of crude oil; middle, how much aggregate-demand shocks explain of the historically observed fluctuations in the real price of crude oil; bottom, how much oil-specific-demand shocks explain of the historically observed fluctuations in the real price of crude oil. Estimates are based on the structural VAR model in (1). The solid, purple line is generated by the original code and data. The dashed, beige line, which lies on top of the purple line, is generated by the replication code and the original data. The dot–dashed, orange line is generated by the replication code using data from February 1974 to December 2007. The dotted, pink line is generated by the replication code using data through March 2024.
The top panel of figure 4 shows that “oil supply shocks historically have made comparatively small contributions to the price of oil” (Kilian, 2009a, 1062).
The middle panel of 4 shows that aggregate-demand shocks cause prolonged and persistent changes in the price of oil. The bottom panel of 4 shows that shocks to oil-specific demand are responsible for the fluctuations in the price of oil. As Kilian (2009a) points out, this statistical understanding of crude-oil prices is inconsistent with narrative hypotheses about supply disruptions in the Middle East. There may be supply disruptions, but the disruptions affect firms that purchase oil, fearing further supply disruptions. These precautionary-demand motives cause the price of oil to fluctuate.
Figure 4 is directly comparable to Kilian’s (2009a) figure 4. The conclusions drawn from the figure largely agree—except for the cumulative effect the aggregate-demand shock has on the real price of crude oil using the updated series through December 2007 (the middle panel). The discrepancy can be explained by the updated series for real global economic activity and the real price of oil in figure 1. The end of the updated series for real activity is characterized by a sharp rise compared to the original series. The beginning of the updated series for the price of oil is lower than the original series. Before 1990 in the updated data, a given level of real economic activity is associated with a lower real price of oil compared to the original data. After 1990 it will not be surprising then that the estimated model using updated data predicts that the real price of oil will be relatively low, producing a negative error, which shows up in the historical decomposition. Comparing the original data to the updated replication series in figure 1 through 2007, the beginning of the series for the real price of oil and the end of the series for real economic activity are challenging for the model to match.
The discrepancy, however, is entirely absent when the structural model is estimated using data through March 2024. Therefore, going forward results will be shown for estimates produced using the full sample.
2.5 The demand and supply shocks in the global oil market affect the US economy in different ways
How does the global market for oil affect economic activity in the United States? Having identified the structural innovations using the model in 1, it is possible to relate the innovations to US statistics like the inflation rate measured by the Consumer Price Index (CPI) and the growth rate of real gross domestic product (GDP).
[…]
$z_{t}=\omega_{j}+\sum_{i=0}^{12}\phi_{ji}\hat{\zeta}_{jt-i}+u_{jt},$ (4)
where $z_{t}$ is either the quarterly growth rate of real GDP or the quarterly inflation rate and $u_{jt}$ is a potentially serially correlated error. As Kilian (2009a) notes, the term $\phi_{jh}$ in the regression model is the impulse response at horizon $h$ of $z$ to structural innovation $j$.
The possibility for serial correlation in (4) is accounted for by constructing confidence intervals using a block-bootstrap procedure. The results presented here use a block length of 6 quarters, whereas Kilian (2009a) uses 4 quarters.
Figure 5 shows how the level of US real GDP and the level of the CPI respond to the structural shocks. The solid, grey lines show impulse responses estimated using the original data and the replication code. The associated two-standard-error bands are shaded in green. The impulse responses estimated using Kilian’s (2009a) code agree exactly, which produces a striped effect. The dotted, blue lines show the upper and lower ends of the confidence intervals using Kilian’s (2009a) code. These estimates produce slightly narrower confidence intervals, likely due to the fact that the block length used to produce the shaded region is two quarters longer. Nevertheless, the difference is minor. Figure 5 is directly comparable to Kilian’s (2009a) figure 5.
Figure 5: Responses of US real GDP and the CPI to identified structural shocks (based on original data)
Note: Clockwise from top left, cumulative response of real GDP to an unanticipated disruption in crude-oil supply; cumulative response of the price level to an unanticipated disruption in crude-oil supply; cumulative response of the price level to an unanticipated increase in aggregate demand; cumulative response of the price level to an unanticipated increase in oil-specific demand; cumulative response of real GDP to an unanticipated increase in oil-specific demand; cumulative response of real GDP to an unanticipated increase in aggregate demand. Solid lines are based on the point estimates from the model in (4). Shaded areas show two-standard-error bands constructed using a block bootstrap procedure. Dotted lines show the point estimates and two-standard error bands produced by Kilian’s (2009a) code. All statistics in the figure are based on the original data (Kilian, 2009b).
Figure 6 updates figure 5. To do this, data from February 1974 to March 2024 are first used to estimate the structural VAR model. Then the associated monthly series of structural shocks are averaged by quarter to produce the quarterly averages used in the second-stage regressions in (4).
Figure 6: Responses of US real GDP and the CPI to identified structural shocks
Note: Clockwise from top left, cumulative response of real GDP to an unanticipated disruption in crude-oil supply; cumulative response of the price level to an unanticipated disruption in crude-oil supply; cumulative response of the price level to an unanticipated increase in aggregate demand; cumulative response of the price level to an unanticipated increase in oil-specific demand; cumulative response of real GDP to an unanticipated increase in oil-specific demand; cumulative response of real GDP to an unanticipated increase in aggregate demand. Solid lines are based on the point estimates from the model in (4). Shaded areas show one- and two-standard-error bands constructed using a block bootstrap procedure. The estimated structural shocks used to produce the responses are based on data from February 1974 to March 2024.
Figure 6 shows that similar patterns emerge when the series for real economic activity is corrected and the latest data are used to construct estimates. An unanticipated oil-supply disruption lowers real GDP on impact. The effect is statistically significant when using the one-standard-error bands for up to nearly three years. The unanticipated oil-supply disruption causes the CPI to increase for nearly 9 quarters. The effect is statistically significant using the one-standard-error bands. Only after 9 quarters does the price level start to decline.
Turning to the second row of 6, an unanticipated expansion in aggregate demand causes real GDP to increase on impact, although the effect is marginally statistically significant. After 5 quarters, real GDP falls below its initial level. The same unanticipated expansion causes the price level to rise for the first 7 quarters after the shock. The response is statistically significant using the one-standard-error bands. This pattern is consistent with the increase in global aggregate demand causing an initial increase in US growth, but the growth is eventually hampered by the higher price of global commodities, including the price of oil, causing real GDP to fall. Kilian (2009a) provides further discussion.
The third row of figure 6 shows that an unanticipated increase in oil-specific demand causes real GDP and the price level to rise. The increase in the price level is highly statistically significant. Eventually the higher prices cause real GDP to fall. Again, Kilian (2009a) provides further discussion.
In summary, we are able to replicate Kilian’s (2009a) substantive results. Patterns uncovered by the historical decomposition and the second-stage impulse response functions hold using data updated through March 2024.
3 Updated responses of oil production, real economic activity, and the real price of oil to demand and supply shocks
Section 2.4 provides evidence that conclusions about the determinants of the real price of crude oil apply through March 2024; namely, disruptions to oil supply contribute little to fluctuations in price, whereas oil-specific demand shocks are responsible for the ups and downs and aggregate-demand shocks cause “long swings” in price (Kilian, 2009a, 1062). Section 2.5 provides evidence that the US economy responds similarly to changes in the oil market over the later period.
To test this stability proposition further, in this section we use updated data to estimate the impulse response functions. In addition, confidence intervals will be produced using the moving-block-bootstrap method introduced by Brüggemann, Jentsch, and Trenkler (2016). Their recursive-design residual-block-bootstrap procedure is robust to conditional heteroskedasticity. In addition, the method is designed to replicate the asymptotic distribution of both the slope parameters and the variance–covariance matrix. In effect, the wild bootstrap “produces confidence intervals that are too narrow and have coverage accuracy that is too low asymptotically” (Kilian and Lütkepohl, 2017, 346–347). An excellent discussion of the issue is provided by Kilian and Lütkepohl (2017). To implement the moving-block bootstrap procedure, a block length of 36 months is chosen.
Figure 7 shows the responses of production, real economic activity, and the real price of oil to one-standard-deviation structural innovations. When comparing figure 7 to figure 3, there are two differences to keep in mind. First, the structural VAR model is estimated using data through March 2024. Second, the confidence intervals are constructed using the procedure introduced by Brüggemann, Jentsch, and Trenkler (2016). Nevertheless, results are remarkably similar. An unexpected disruption to oil supply causes an immediate decline in oil production and the decline persists. The unexpected fall in oil production initially has no effect on bulk-dry-cargo shipping rates relative to trend, but the measure of real economic activity picks up after 4 months. The rise could reflect higher fuel costs or expected higher fuel costs (relative to the US consumer price index, which is used to deflate the series). For example, the next panel shows that the unexpected disruption to supply causes an even more pronounced rise in the price of oil. The rise in the price of oil is statistically significant at all horizons based on the one-standard-error bands.
Figure 7: Responses to one-standard-deviation structural shocks
Note: Solid lines show the point estimates and the shaded regions show one- and two-standard-error bands. Estimates are based on the structural VAR model in (1), using the updated data (February 1974–March 2024). Confidence intervals were constructed using a residual block bootstrap procedure (Brüggemann, Jentsch, and Trenkler, 2016).
Turning to the second row of figure 7, over the entire sample, oil production seems to respond more acutely to an unanticipated increase in aggregate demand, although a months-long delay remains. The unanticipated rise in aggregate demand causes a large, significant jump in global real economic activity, but the effect is less persistent compared to the effect estimated over only the earlier period. And, compared to only the earlier period, there is a more pronounced rise in the real price of oil in response to an unanticipated increase in aggregate demand. Nevertheless, the patterns are overwhelmingly consistent with Kilian’s (2009a) substantive conclusions.
The bottom row of figure 7 shows that an unanticipated rise in oil-specific demand causes an immediate rise in the real price of oil. Again there is evidence of a precautionary-demand motive. Like over the earlier period, the unanticipated rise in oil-specific demand is also associated with an increase in real economic activity and a rise in oil production. Perhaps, unlike in the earlier period, oil production responds more to surprises in oil-specific demand, although there is a delay of two months and the effect is significant using the one-standard-error bands only.
In summary, the substantive conclusions of Kilian (2009a) hold when all available data are used and confidence sets are constructed using a method robust to potential conditional heteroskedasticity.
4 Application to the clean-energy transition
Much previous interest has centered around the role that oil-supply and -demand shocks have on macroeconomic outcomes like inflation, including responses by the Federal Reserve (see Bernanke, Gertler, and Watson, 1997, for a prominent example). Another compelling reason to study these types of shocks is understanding how local economies depend on oil.
Certain local economies may be especially hard hit as the US transitions away from fossil fuels to sources of clean energy. Previous structural transitions—such as the transition away from coal and the transition to work in the United States being replaced by machines and jobs elsewhere—are causes for concern. In California, policymakers have adopted the goal of carbon neutrality by 2045 or earlier (CARB, 2022). Kern County produces 70 percent of all oil produced in California. As California transitions away from oil, major concerns for Kern include lost jobs and lost tax revenue.
On such a granular level, it is nearly impossible to determine how much of Kern’s economy depends on oil. Yet, one way to understand the linkages between Kern and the global market for oil is by assessing how Kern’s unemployment rate responds to structural innovations identified by the statistical model in (1), using second-stage regressions like those in (4). Specifically, we regress the seasonally adjusted, monthly unemployment rate in Kern on lags of the estimated structural shocks:
$\text{unemployment rate in Kern}_{t}=\kappa+\sum_{i=0}^{12}\psi_{ji}\hat{\varepsilon}_{jt-i},$ (5)
where $\hat{\varepsilon}_{jt}$ refers to the estimated residual in (1) for the $j$th structural shock in month $t$. The impulse response coefficients $\psi_{jh}$ are the response of Kern’s unemployment rate to structural-innovation $j$ at horizon $h$. Inference is based on a block-bootstrap procedure where the block length is chosen to be half a year.
Data on the unemployment rate in Kern County, California, are available from the Local Area Unemployment Statistics program of the Bureau of Labor Statistics starting in January 1990. The data extend through March 2024.
Figure 8 shows the estimated coefficients. The solid, beige lines show the impulse-response coefficients. The shaded areas show the one- and two-standard-error bands.
Figure 8: Responses of the unemployment rate in Kern County, California, to identified structural shocks
Note: Top, response of the unemployment rate to an unanticipated disruption in oil supply; middle, response of the unemployment rate to an unanticipated increase in aggregate demand; bottom, response of the unemployment rate to an unanticipated increase in oil-specific demand. Solid lines show the point estimates from the model in (5). Shaded areas show one- and two-standard-error bands constructed using block bootstrap procedure.
The top panel shows that, upon impact, an unanticipated disruption in oil production has little effect on the unemployment rate in Kern. And there is some evidence that the unemployment rate gradually moves upward, but the evidence is not statistically significant. The response is not obvious, as a disruption to global supply from Kern’s perspective could have represented a windfall increase in the price of oil, which would have benefited the local economy. This does not appear to be the case. Unemployment rises on impact and remains higher a year after the initial disruption, although the effect is marginally statistically significant.
In contrast, an unanticipated one-standard-deviation increase in aggregate demand causes unemployment to decline by over 1 percentage point 4 months after the shock. By 8 months, the unemployment rate falls by 2 percentage points. A year after the shock, the unemployment rate is 3 percentage points lower. The decline is statistically significant at all horizons.
The bottom panel shows that an unanticipated increase in oil-specific demand is also good for the Kern economy. A year after the initial increase the unemployment rate is 2.5 percentage points lower, having fallen throughout the year after the shock. The effect is significant based on the one-standard-error band.
These effects are meaningful. The expanded economic opportunities after an increase in oil-specific demand suggests that Kern may well depend on oil in a meaningful way and the transition away from fossil fuels may be a challenge. Further research on this topic will be aided by having the option to work within the R ecosystem.
Figure 8 makes clear, though, that Kilian’s (2009a) point about identifying the origin of the shock is paramount. Even though the price of oil may be determined independently of economic activity in Kern County, the dynamic effect of a negative supply disruption differs from the dynamic effect of a positive demand innovation—the first causes the unemployment rate to rise, while the second causes the unemployment rate to fall.
5 Conclusion
In this paper we successfully replicated the substantive conclusions made by Kilian (2009a); namely, not all oil-price shocks are alike. Understanding the origins of such shocks is necessary for policymakers. For example, as Kilian (2009a) points out, a positive structural innovation to the global business cycle will cause national income in the United States to increase—as the United States directly benefits from an increase in global economic activity—but the expansion of US income will be hampered by the rising cost of oil that inevitably follows. A negative innovation to crude-oil production will cause national income in the US to fall and cause the CPI to rise. Both innovations initially cause the price of crude oil to rise, but the dynamic effects differ meaningfully. Policymakers will need to understand the origin of the shock in order to respond appropriately.
Kilian (2009a) should be commended for his excellent code. The original replication package reproduces the results even when we port the code to R (Kilian, 2009b). In fact, the original replication package offers an excellent starting point for further analysis.
We furthered the analysis in several ways:
  • We ported the code to R, which will benefit researchers interested in working within the R ecosystem. For many, tidying data in R is much easier than tidying data in MATLAB.
  • Our replication and paper are written using the targets package. We demonstrate how to use this Make-like took to ensure reproducibility. The targets package can also be used to exploit parallel computing, which could be used to speed the construction of bootstrapped confidence intervals.
  • Data from the US Energy Information Administration and FRED database are retrieved by code. This makes updating the structural shocks straightforward.
  • We provided an application of Kilian’s (2009a) procedure for identifying structural shocks. These shocks are used learn about how a local labor market responds to the global oil market. Understanding linkages like these may well be more important as society transitions toward sustainable energy sources.